# Video-XL-Pro: Reconstructive Token Compression for Extremely Long Video Understanding

### v1

@@ 3.1 Overview
The overview of Video-XL-Pro is divided into two stages: self-supervised pertaining for reconstructive token compression and MLLM training, as illustrated in Figure 2. In the first stage, our goal is to develop a robust visual encoder capable of generating compact and comprehensive visual tokens using self-supervised learning methods. In Section 3.2, we provide a detailed explanation of the two key components: Dynamic Token Synthesizer (DTS) and Semantic-Guided Masking (SGM). In the second stage, we utilize pretrained SigLIP and DTS as our visual encoder and follow Video-XL for fine-tuning. Additionally, we introduce a video dataset pruning strategy and a Query-Aware Selector to improve training efficiency, as detailed in Section 3.3.
@@ 3.2 Reconstructive Token Compression
Model Size Training Data (M) MLVU Dev MLVU Test VideoMME VNBench LongVideo MVBench
M-avg G-avg M-avg G-avg W/o sub W sub
@@
LongVU-3.2B [36] 3B 3.2/0.5 55.9 - - - 51.5 - - - -
InternVL2.5-2B [5] 2B - 58.9 - - - 51.9 54.1 - 52.0 68.8
Apollo-3B [57] 3B 0.2/0.4/3.2 65.0 - - - 58.4 60.6 - - -
VideoChat-Flash-2B [18] 2B - 65.7 - - - 57.0 63.9 - 58.3 70.0
VideoLLaMA3-2B [49] 2B - 65.4 - - - 59.6 63.4 - 57.1 65.5
@@
LongVA-7B [51] 7B - 56.3 4.33 41.1 3.91 52.6 54.3 41.5 47.8 -
Pixtral-12B [1] 7B - - - - - 40.7 47.5 - - -
LongVU-7B [36] 7B 3.2/0.5 65.4 - - - 60.6 - - - -
LLaVA-OV-7B [15] 7B - 64.7 - - - 58.2 61.5 - - -
Video-XL [37] 7B 2/1 64.9 4.50 45.5 4.21 55.5 61.0 61.6 50.7 55.3
TimeMaker-8B [4] 8B 45 - - - - 57.3 62.8 - - -
Kangaroo-8B [22] 8B 372 61.0 - - - 56.0 57.6 - 54.8 -
Apollo-7B [57] 7B 0.2/0.4/3.2 67.3 - - - 61.3 63.3 - - -
Video-XL-Pro 3B 2.0/0.8 68.1 5.27 46.9 4.51 57.9 62.6 63.4 51.5 57.6
@@ 3.3 MLLM Training
Following the success of the MLLM paradigm, which leverages visual instruction tuning to align vision and language representations, Video-XL-Pro consists of three main modules: the LLM backbone, Vision Encoder, and Cross-Modality Projector. We select Qwen-2.5 [47] with 3B parameters as our LLM, due to its strong language understanding and reasoning capabilities. Following Video-XL [37], we apply KV sparsification, compressing tokens in the LLM by a factor of 16×. For the Vision Encoder, we utilize the pretrained visual encoder, which consists of SigLIP and DTS, enabling the generation of comprehensive and compact visual tokens. Following LLaVA [21], we employ a two-layer MLP with a GELU activation function as a Cross-Modality Projector to align the Vision Encoder and LLM. Besides, we propose the Query-aware selector for selecting informative tokens that permanent to the given query, and propose video dataset pruning for improving efficiency.
@@ 4.1 Training Data
We use large-scale image-text and video-text datasets for pre-training, aligning vision and language modalities. For image-text pairs, we collect 2M samples, removing low-quality data with noisy labels while retaining those with detailed captions. For video-text pairs, we apply video-length-based filtering, retaining shorter videos with minimal scene changes. Ultimately, we collect 2M image samples and 400K video samples from LLaVA-OneVision [15], LAION-2M [11], OpenViD [28], and Charades [38]. In the fine-tuning stage, we follow the Video-XL settings, utilizing Bunny-695K [11] image data and our pruned 550K video data from LLaVA-178K [53], VCGPlus [25] for training.
@@ 4.2 Implementation Details
In the pretraining stage, only the projector is trained, with a learning rate of 5e-4. In the fine-tuning stage, all parameters are trained, with the learning rate set to 1e-5. The batch sizes for pretraining and fine-tuning are 8 and 1, respectively. We employ a cosine learning rate schedule with a warm-up ratio of 0.03 in both stages. In the fine-tuning stage, videos are sampled at 1 FPS, and if the video length exceeds 360 frames, we uniformly sample 360 frames. All experiments are conducted on 32×32$\times$32 × A800-80GB GPUs.
@@ 4.3 Benchmarks and Metrics
We empirically evaluate the effectiveness of Video-XL-Pro based on several popular long video understanding benchmarks. 1. MLVU [55], a comprehensive benchmark which is made up of both multiple choice and generation tasks. 2. Video-MME [7], another extensive benchmark covering videos of diverse genres and lengths (short, medium, and long). 3. VNBench [54], a synthetic benchmark focused on assessing models’ ability to handle long-video tasks, such as retrieval, ordering, and counting. 4. LongVideoBench [45], a benchmark designed for tasks that require precise retrieval and reasoning over detailed multi-modal information within extended inputs.
@@ 4.4 Main Results
We present the performance of Video-XL-Pro on popular long video benchmarks in Table 1. Notably, it outperforms existing methods on both the Dev and Test set of MLVU. Despite having only 3B parameters, it even surpasses GPT-4o on MLVU and outperforms the majority of existing 7B models, including Apollo 7B, which was trained on 5M data, by 0.8%, demonstrating highly competitive results. For Video-MME, Video-XL-Pro achieves accuracies of 57.9% and 62.6% under the “no subtitles” and “with subtitles” settings, respectively, reaching state-of-the-art levels compared to advanced models. In VNBench, Video-XL performs the best among open-source models, leading the previous best model by nearly 10%. In this evaluation, it once again surpasses GPT-4V and achieves performance comparable to GPT-4o. Last but not least, although Video-XL-Pro is primarily designed for long video understanding tasks, it also excels in short video tasks, achieving competitive results on the MVBench benchmark.
@@ 4.7 Ablation Studies
SigLIP+DTS 576 98.17/89.22 81.07 99.33
@@
SigLIP+DTS 144 95.49 67.82 68.7
Table 2: Effectiveness of Dynamic Token Synthesize Method.
@@
SigLIP+DTS 144 53.1 47.5 1288.7 47.2
@@ 4.8 Qualitative Results
We employ qualitative evaluation to intuitively analyze Video-XL-Pro. In this experiment, we compare it with longvu and Qwen2.5-VL 7B based on ultra-long videos (with durations exceeding 30 minutes). As shown in Figure 6, Video-XL-Pro is capable of accurately locating key information in long videos and presenting its details; in contrast, Qwen2.5-VL struggles with understanding the content of long videos and fine-grained details. Video-XL-Pro effectively locates the required key information in long videos. This further validates the robustness of Video-XL-Pro in long video understanding tasks.
@@ Appendix A Training Datasets
Video 30K CinePine
Video 40K Sharegpt4v
@@
Video 4K ScienceQA
Video 100K VCG
Video 16K ActivitynetQA
Image 480K Bunny695k

### v2

@@ 3.1 Overview
The overview of Video-XL-Pro is divided into two stages: self-supervised pertaining for reconstructive token compression and MLLM training, as illustrated in Figure 2. In the first stage, our goal is to develop a robust visual encoder capable of generating compact and comprehensive visual tokens using self-supervised learning methods. In Section 3.2, we provide a detailed explanation of the two key components: Dynamic Token Synthesizer (DTS) and Semantic-Guided Masking (SGM). In the second stage, we utilize pretrained SigLIP and DTS as our visual encoder and follow VLM for fine-tuning. Additionally, we introduce a video dataset pruning strategy to enhance training efficiency, along with a Query-Aware Selector to expand the maximum input context length supported by VIdeo-XL-Pro, as detailed in Section 3.3.
@@ 3.2 Reconstructive Token Compression
Model Size Training Data (M) MLVU Dev MLVU Test VideoMME VNBench LongVideo TempCompass
W/o sub W sub
@@
InternVL2.5-2B [6] 2B - 61.4 - 51.9 54.1 - 52.0 -
InternVL2.5-4B [6] 4B - 68.3 - 62.3 63.6 - 55.2 -
VideoChat-Flash-2B [23] 2B - 65.7 - 57.0 63.9 - 58.3 -
VideoLLaMA3-2B [58] 2B 15.6M/22M/19M/5.7M 65.4 - 59.6 63.4 - 57.1 63.4
LongVU-3.2B [43] 3B 3.2M/0.5M 55.9 - 51.5 - - - -
Apollo-3B [66] 3B 0.2M/0.4M/3.2M 65.0 - 58.4 60.6 - 55.1 62.5
Qwen2.5-VL-3B [3] 3B - 68.2 - 61.5 67.6 - 54.2 64.4
@@
LongVA-7B [60] 7B - 56.3 41.1 52.6 54.3 41.5 47.8 57.0
Pixtral-12B [1] 7B - - - 40.7 47.5 - - -
LongVU-7B [43] 7B 3.2M/0.5M 65.4 - 60.6 - - - -
Oryx1.5 [29] 7B - 63.8 - 59.0 - - - -
LLaVA-OV-7B [20] 7B - 64.7 - 58.2 61.5 - - -
Video-XL [44] 7B 2.0M/1.0M 64.9 45.5 55.5 61.0 61.6 50.7 -
TimeMaker-8B [5] 8B 45M - - 57.3 62.8 - - 60.4
Apollo-7B [66] 7B 0.2M/0.4M/3.2M 67.3 - 61.3 63.3 - 58.5 64.9
Video-XL-Pro 3B 2.0M/1.0M 70.6 47.0 60.0 64.3 63.0 56.7 65.7
@@
Method Size V-STaR
$R@1_{(\text{IOU}=0.3)}$ $R@1_{(\text{IOU}=0.5)}$ $R@1_{(\text{IOU}=0.7)}$ mIoU
VTimeLLM [16] 7B 25.24 10.88 3.15 17.13
TimeChat [41] 7B 17.80 8.68 3.48 12.01
TRACE[10] 7B 28.53 14.17 6.73 19.74
InternVL-2.5 [6] 8B 11.98 4.87 2.34 8.72
Qwen2.5-VL-7B [3] 7B 17.03 8.92 3.72 11.48
Video-LLaMA3 [58] 7B 35.73 19.80 8.68 22.97
Video-XL-Pro 3B 40.72 25.78 15.58 25.07
Table 2: Results on temporal grounding task, Performance Comparison on V-STaR.
Method Size Charades-STA
$R@1_{(\text{IOU}=0.3)}$ $R@1_{(\text{IOU}=0.5)}$ $R@1_{(\text{IOU}=0.7)}$ mIoU
VTimeLLM [16] 7B 51.0 27.5 11.4 31.2
TimeChat [41] 7B - 32.2 13.4 -
TimeSuite [56] 7B 69.9 48.7 24.0 -
Qwen2.5-VL-7B [3] 7B - - - 43.6
VideoChat-Flash [23] 7B 72.5 51.4 26.4 48.0
Video-XL-Pro 3B 80.3 64.1 42.7 51.8
Table 3: Results on temporal grounding task, Performance Comparison on Charades-STA.
@@ 3.3 MLLM Training
Following the success of the MLLM paradigm, which leverages visual instruction tuning to align vision and language representations, Video-XL-Pro consists of three main modules: the LLM backbone, Vision Encoder, and Cross-Modality Projector. We select Qwen-2.5 [55] with 3B parameters as our LLM, due to its strong language understanding and reasoning capabilities. For the Vision Encoder, we utilize the pretrained visual encoder, which consists of SigLIP and DTS, enabling the generation of comprehensive and compact visual tokens. Following LLaVA [26], we employ a two-layer MLP with a GELU activation function as a Cross-Modality Projector to align the Vision Encoder and LLM. Besides, we propose the Query-aware selector for selecting informative tokens that permanent to the given query, and propose video dataset pruning for improving efficiency.
@@ 4.1 Training Data
We use large-scale image-text and video-text datasets for pre-training, aligning vision and language modalities. For image-text pairs, we collect 2M samples, removing low-quality data with noisy labels while retaining those with detailed captions. For video-text pairs, we apply video-length-based filtering, retaining shorter videos with minimal scene changes. Ultimately, we collect 2M image samples and 400K video samples from LLaVA-OneVision [20], LAION-2M [14], OpenViD [34], and Charades [45]. In the fine-tuning stage, the following datasets are collected: Bunny [14], Sharegpt-4o [15], and OCR-VQA [33], These datasets are combined with our pruned video data, which contains NExT-QA [53], Sharegpt-4o [15], CinePile [39], LLaVA-178K [62], VCGPlus [31] for training.
@@ 4.2 Implementation Details
In the pretraining stage, only the projector is trained, with a learning rate of 5e-4. In the fine-tuning stage, all parameters are trained, with the learning rate set to 1e-5. The batch sizes for pretraining and fine-tuning are 8 and 1, respectively. We employ a cosine learning rate schedule with a warm-up ratio of 0.03 in both stages. In the fine-tuning stage, videos are sampled at 1 FPS, and if the video length exceeds 240 frames, we uniformly sample 240 frames. All experiments are conducted on 32×32$\times$32 × A800-80GB GPUs.
@@ 4.3 Benchmarks and Metrics
We empirically evaluate the effectiveness of Video-XL-Pro based on several popular long video understanding benchmarks. 1. MLVU [64], a comprehensive benchmark that consists of multiple choice and generation tasks. 2. Video-MME [9], another extensive benchmark covering videos of diverse genres and lengths (short, medium, and long). 3. VNBench [63], a synthetic benchmark focused on assessing models’ ability to handle video tasks, such as retrieval, ordering, and counting. 4. LongVideoBench [52], a benchmark designed for tasks that require precise retrieval and reasoning over detailed multi-modal information within extended inputs. 5. Tempcompass [28], a benchmark designed for evaluating visual-language models’ temporal understanding of long videos. We further evaluate the model’s timestamp awareness using the Charades-STA [19] temporal grounding dataset and V-STaR [7] long video temporal grounding dataset.
@@ 4.4 Main Results
We present the performance of Video-XL-Pro on popular long video benchmarks in Table 1. Notably, it outperforms existing methods on both the Dev and Test set of MLVU. Despite having only 3B parameters, its performance on MLVU even surpasses GPT-4o, as well as well-known open-source models with similar parameter scales, such as Qwen2.5-VL-3B and InternVL2.5-4B. It also outperforms the majority of existing 7B models, including Apollo 7B, which was trained on 3.2M sft data, by 3.3%, demonstrating highly competitive results. For Video-MME, Video-XL-Pro achieves accuracies of 60.0% and 64.3% under the “no subtitles” and “with subtitles” settings, respectively, reaching state-of-the-art levels compared to advanced models. In TempCompass, Video-XL-Pro achieves the best performance among open-source models, even outperforming Qwen2.5-VL-3B by 1.3% and Apollo 7B by 0.8%, especially considering it is trained with far less data. It also achieves competitive results on LongVideoBench. Last but not least, although Video-XL-Pro is primarily designed for long video understanding tasks, it also excels in short video tasks, achieving competitive results on the VNBench benchmark.
@@
4.5 Temporal Grounding
As detailed in Table 2 and Table 3, compared to previous methods that rely on complex architectures to perceive timestamps, our method demonstrates surprisingly strong temporal grounding capabilities, especially on the long video temporal grounding benchmark V-STaR, where it almost outperforms the existing 7B MLLMs and achieves highly competitive performance.
@@ 4.8 Ablation Studies
To demonstrate the contribution of Semantic-Guided Masking, we fixed the compression rate of DTS and compared it with a random mask at the same ratio. As shown in Table  6, DTS with Semantic-Guided Masking achieved improvements on both image and video understanding benchmarks. This indicates that Semantic-Guided Masking can enhance the important features in visual understanding for DTS, thereby strengthening the model’s comprehension ability.
@@
SigLIP+DTS (2×2$\times$2 ×) 576 98.07/88.65 80.59 98.91
SigLIP+DTS (4×4$\times$4 ×) 576 98.17/89.22 81.07 99.33
SigLIP+DTS (8×8$\times$8 ×) 576 98.07/88.65 80.59 98.91
@@
SigLIP+DTS (2×2$\times$2 ×) 256 93.97 67.56 68.3
SigLIP+DTS (4×4$\times$4 ×) 144 95.49 67.82 68.7
SigLIP+DTS (8×8$\times$8 ×) 72 89.27 65.94 66.4
Table 4: Effectiveness of Reconstructive Token Compression Method.
@@
SigLIP+DTS (2×2$\times$2 ×) 256 52.0 46.7 1256.7 46.7
SigLIP+DTS (4×4$\times$4 ×) 144 53.1 47.5 1288.7 47.2
SigLIP+DTS (8×8$\times$8 ×) 72 46.2 43.1 1224.9 45.3
@@
Model MLVU VideoMME MME MMB
SigLIP+DTS+RandomMask 52.4 47.0 1260.4 46.1
SigLIP+DTS+SGM 53.1 47.5 1288.7 47.2
Table 6: Analysis of Semantic-Guided Masking in ReCot.
@@ 4.9 Qualitative Results
We employ qualitative evaluation to intuitively analyze Video-XL-Pro. In this experiment, we compare it with Qwen2.5-VL-7B based on ultra-long videos (with durations exceeding 30 minutes). As shown in Figure 6, Video-XL-Pro is capable of accurately locating key information in long videos and presenting its details; in contrast, Qwen2.5-VL struggles with understanding the content of long videos and fine-grained details. Video-XL-Pro effectively locates the required key information in long videos. This further validates the robustness of Video-XL-Pro in long video understanding tasks.
@@ Appendix A Training Datasets
Video 28K CinePine
Video 38K Sharegpt4v
@@
Video 3K ScienceQA
Video 107K VCG
Video 15K ActivitynetQA
Image 390K Bunny695k